OpenStudy (anonymous):

how do you divide 1/4 and 1/3

OpenStudy (anonymous):

do you mean dividing them with each other?

OpenStudy (anonymous):

1 1 1 3 3 --- divided by --- = -----*----- = ---- 4 3 4 1 4

OpenStudy (saifoo.khan):

OMG!

OpenStudy (anonymous):

your teacher wants to make a smaller serving. she is going to only make 1/3 times as much. how much will she need? 1/4 cup, 1/3 cup, 2/5 cup, 1/6cup, 3/8 cup, 2/7 cup, 2/9 cup. each one needs and answer. for the first one he got 1/4 and 1/3 + 1/12

OpenStudy (saifoo.khan):

\[\Large \frac{\frac14}{\frac13}\]\[\Large \frac14 \div \frac13\]\[\Large \frac14 \times \frac31\] \[\Large \frac34\]

OpenStudy (anonymous):

when u hv to divide a fraction by another fraction, you change division to multiplication and take the multiplicative inverse of the second fraction as I hv done...

OpenStudy (saifoo.khan):

Can u please rewrite the question only? and is that question compelete?

OpenStudy (saifoo.khan):

Complete*

OpenStudy (saifoo.khan):

which problem are you talking about?

OpenStudy (anonymous):

recipe 1/4 cup of tabasco, 1/3 cup, 2/5 cup, 1/6 cup, 3/8 cup, 2/7 cup, 2/9 cup. your teacher wants to make a smaller serving. she is going to make only 1/3 times as much. How much will she need of each ingrdient?

OpenStudy (saifoo.khan):

what are: 1/3 cup, 2/5 cup, 1/6 cup, 3/8 cup, 2/7 cup, 2/9 cup. ?

OpenStudy (saifoo.khan):

Ok, well as much i get the question, i think, we will add all the fractions first, then divide the result with 3.

OpenStudy (saifoo.khan):

Can u post the photo of the question?

OpenStudy (saifoo.khan):

umm.. i got the question but i dont get what are other fractions for? 1/3 cup, 2/5 cup, 1/6 cup, 3/8 cup, 2/7 cup, 2/9 cup.

OpenStudy (saifoo.khan):

recipe for?

OpenStudy (saifoo.khan):

i mean like if u are cooking something, it's like, 3/4 cups of floor. 1/2 tea spoon of salt 3/2 tea spoon of spice. so are these like those in the question?

OpenStudy (saifoo.khan):

Go for it, try adding all the fractions, then divide the result by 3.

OpenStudy (saifoo.khan):

Sure! :)

OpenStudy (saifoo.khan):

Did u saved it right? Now click the "attach file" and locate the file where you saved it. select the file, click ok. now wait for a bit. and then click post.

OpenStudy (anonymous):

where is repost

OpenStudy (saifoo.khan):

Got it , let me check.. :)

OpenStudy (saifoo.khan):

Welcome..

OpenStudy (saifoo.khan):

Perfect. Same thing as i figured. Add the fractions on top. Then multiply the result by: 1) 1/3 2) 1/2 3) 5/6 4) 1/4

OpenStudy (saifoo.khan):

umm. hold a sec please let me check. :)

OpenStudy (saifoo.khan):

OH! im really sorry i did a mistake!

OpenStudy (saifoo.khan):

i made the question harder. :( sowwie.

OpenStudy (saifoo.khan):

Sure, okay.. so.. let's do the 4rd part. \[Tasbasco sause = \frac14 \times \frac56\] \[Red onions = \frac13 \times \frac56\]and so on.

OpenStudy (anonymous):

how was question 1, 2, 3 ,4

OpenStudy (saifoo.khan):

Sorry again, i showed working for 3rd part. typos! >_<

OpenStudy (saifoo.khan):

For forth question, sause = 1/4 * 1/4 onions = 1/3 * 1/4 and so on

OpenStudy (anonymous):

do we reduce

OpenStudy (anonymous):

* multiply

OpenStudy (saifoo.khan):

\[\frac14 \times 14 = \frac{1}{16}\] \[\frac13 \times \frac 14 = \frac{1}{12}\]

OpenStudy (saifoo.khan):

Yes ma'am.

OpenStudy (anonymous):

so all of them are just multiplying and reducing

OpenStudy (saifoo.khan):

YES!

OpenStudy (anonymous):

Thank you so much for your time and being so patient with a mom now i can explain it. how do i delete all my old questions

OpenStudy (saifoo.khan):

No problem! And And And Please dont forget to spread the Openstudy word. Delete questions why?

OpenStudy (anonymous):

i have so many repeats of the same ones

OpenStudy (anonymous):

i will spread it everyone is so kind

OpenStudy (saifoo.khan):

Cheers! :) Be back soon!